OpenStudy (anonymous):

(x+3)(x^4-4x^2+5) can you show me the steps to to simplify

OpenStudy (sasogeek):

ok so here what we have to do is simplify :P obviously but the question is how... right?

OpenStudy (anonymous):

i just stared to learn this and i need to know what steps to do first

OpenStudy (sasogeek):

well i want to engage you so that you can understand it better :)

OpenStudy (anonymous):

ok so what step do i do first

OpenStudy (sasogeek):

ok so with simplification of such an expression, you just have to use the distributive property :) do you know what that is?

OpenStudy (sasogeek):

well this is what the distributive property says, \[\huge a(b+c)=ab+bc \] but in this case, we have something a little bit different so what we have to do is expand this \[\huge (x+3)(x^4-4x^2+5)\] which will become \[\huge x(x^4-4x^2+5)+3(x^4-4x^2+5) \] do you see what i did there?

OpenStudy (phi):

Does the question ask you "to simplify" ? Because all you can do is either (1) expand or (2) factor. Neither is really simplifying in this case.

OpenStudy (sasogeek):

well it is already factored...

OpenStudy (anonymous):

\[x^2=t\] \[t^2-4t+5=t^2-4t+4+1=0\] \[(t-2)^2+1=0\] \[t-2=\pm i\] \[t=2\pm i\] \[x^2=2\pm i\] \[x=\pm \sqrt{2\pm i}\]

OpenStudy (anonymous):

\[(x-3)(x-\sqrt{2-i})(x-\sqrt{2+i})(x+\sqrt{2-i})(x+\sqrt{2+i})\]